### v1

# Activation Steering for Robust Type Prediction in CodeLLMs

Abstract
Contemporary LLMs pretrained on code are capable of succeeding at a wide variety of programming tasks. However, their performance is very sensitive to syntactic features, such as the names of variables and types, the structure of code, and presence of type hints. We contribute an inference-time technique to make CodeLLMs more robust to syntactic distractors that are semantically irrelevant. Our methodology relies on activation steering, which involves editing internal model activations to steer the model towards the correct prediction. We contribute a novel way to construct steering vectors by taking inspiration from mutation testing, which constructs minimal semantics-breaking code edits. In contrast, we construct steering vectors from semantics-preserving code edits. We apply our approach to the task of type prediction for the gradually typed languages Python and TypeScript. This approach corrects up to 90% of type mispredictions. Finally, we show that steering vectors calculated from Python activations reliably correct type mispredictions in TypeScript, and vice versa. This result suggests that LLMs may be learning to transfer knowledge of types across programming languages.
1 Introduction
Large Language Models trained on code (CodeLLMs) are one of the most successful applications of LLMs to date. They are the foundation for both developer tools (e.g., Weiss & Yahav (2013); git (2021)) and for reasoning agents (e.g., Zelikman et al. (2024); Yang et al. (2024); Hu et al. (2024); Li et al. (2022)). However, despite their utility, CodeLLMs can still be unreliable. Recent work shows that even the most capable models are not robust to small semantics-preserving changes such as refactorings and variable renamings (Hooda et al., 2024; Tambon et al., 2024). Thus CodeLLMs often cannot be deployed without careful supervision, such as a human-in-the-loop or robust code sandboxing.
In this paper, we present an approach to make CodeLLMs more robust using activation steering (Li et al., 2024; Rimsky et al., 2023; Turner et al., 2023; Subramani et al., 2022). Activation steering is an inference-time model editing technique that modifies the intermediate computations (activations) of a model using steering vectors to steer model behavior toward desired outcomes. We present a novel technique for constructing steering vectors for the code domain by taking inspiration from mutation testing (DeMillo et al., 1978). Mutation testing automatically constructs minimal semantics-breaking code edits. In contrast, we construct steering vectors from minimal semantics-preserving code edits that lead to CodeLLM mispredictions. The nature of code allows us to construct these code edits in a sound and scaleable way.
[…]
Finally, we are surprised to find that type prediction steering transfers across programming languages. In particular, the steering vectors that we construct from TypeScript prompts are very effective at steering Python type predictions, and vice versa. We conduct a careful analysis and conclude that is it likely that multi-lingual CodeLLMs are learning a representation of types that is shared across the programming languages we study (§ 4.5).
2 Background and Related Work
return s[::-1] == s
   return s[::-1] == s<fim_middle>
Figure 1: An example of a code-infilling prompt using FIM special tokens. The correct prediction expected after the FIM middle token is str.
CodeLLMs and fill-in-the-middle
CodeLLMs are capable of performing a variety of downstream tasks such as generating code from natural language, explaining code, generating tests, and more (Nam et al., 2024; Schäfer et al., 2024). Contemporary CodeLLMs are decoder-only transformers trained on vast amounts of source code (Li et al., 2023; Rozière et al., 2024; Guo et al., 2024). The aforementioned models are also trained to fill-in-the-middle (FIM) (Bavarian et al., 2022; Fried et al., 2023). FIM training does the following (1) It splits $\approx 50\%$ of training items into three chunks—prefix, middle, and suffix—of random lengths; (2) It adds special tokens to the start of each chunk to demarcate chunk boundaries; and (3) It reorders the middle chunk to appear last. The language modelling training objective remains unchanged. Thus FIM inference allows us to generate the middle chunk, conditioned on the prefix and the suffix (figure 1). The resulting model can still be used for conventional left-to-right generation.
Neural type prediction
In this paper we study activation steering for type prediction, formulated as a fill-in-the-middle task. Prior work uses this formulation to evaluate base models (Yee & Guha, 2023; Fried et al., 2023), but does not study models’ internal mechanisms or try to improve task performance in a targeted way. There is prior work that trains smaller, specialized models for type prediction (Hellendoorn et al., 2018; Jesse et al., 2022; 2021; Pandi et al., 2021; Wei et al., 2020), but contemporary CodeLLMs outperform these specialized models (Yee & Guha, 2023; Fried et al., 2023).
Classical type prediction and type inference
Type prediction is also known as type migration, and is distinct from type inference as found in languages such as OCaml and Haskell. In those languages, every variable is typed, even if the types are implicit (Harper & Mitchell, 1993). In contrast, gradual type prediction can change program semantics (Phipps-Costin et al., 2021). There is work on gradual type prediction that uses human-written constraints and constraint solvers (Phipps-Costin et al., 2021; Rastogi et al., 2012; Siek & Vachharajani, 2008; Campora et al., 2018; Henglein & Rehof, 1995; Cartwright & Fagan, 1991). But, these papers present algorithms for variations of the lambda calculus or simple functional languages such as Scheme, and have not been scaled to more complex, modern programming languages.
Mechanistic interpretability
Interpretability seeks to understand the causal mechanisms behind model inference. In particular, mechanistic interpretability studies how model behaviors arise from specific components within the transformer architecture. One of the driving motivations of the field is that by understanding how a model works, we can understand how to align models to desired behavior.
Prior work in mechanistic interpretability includes localizing and editing factual associations within the transformer (Meng et al., 2022), as well as probing hidden representations for knowledge of high-level concepts (Li et al., 2024; Dong et al., 2023). These works are forms of implicit evaluation of model ability, as opposed to explicit evaluation using benchmarks (Dong et al., 2023). Mechanistic interpretability has contributed a wealth of methods for implicit evaluation, for example activation patching (Vig et al., 2020; Variengien & Winsor, 2023), which applies patches to model activations in order to produce a change in behavior. Research leveraging activation patching has suggested the existence of task vectors (Hendel et al., 2023; Ilharco et al., 2022). Task vectors are representations within the transformer which encode a high-level notion of the task described by the prompt. The existence of task vectors is further supported by work in activation steering. Steering has been employed to diminish model deceitfulness and sycophancy (Rimsky et al., 2023; Li et al., 2024), which suggests that steering works by modifying the activations corresponding to task vectors. We thus draw on the idea of task vectors to perform our steering experiments. Importantly, steering relies on the linear representation hypothesis, which argues that concepts are represented as directions in the embedding space of a model (Park et al., 2023). We posit that successful steering performs transformations over model activations that shift the task representation towards a target direction.
Despite much work on the mechanistic interpretability of LLMs, similar research on CodeLLMs has been limited. In this paper we focus on types, which are a fundamental feature of programming languages. One obstacle in this domain is that within a program, a type annotation is semantically constrained by both the tokens that precede it and those that succeed it. For example, the type of a function argument is constrained by how the argument is subsequently used in the function body. To overcome this, we use CodeLLMs that are trained on FIM. This allows us to construct prompts for type prediction.
Mutation testing and program transformations
We take inspiration from mutation testing (DeMillo et al., 1978), but our technique is different. The goal of mutation testing is to test a program’s test suite. To do so, a mutator injects small bugs that alter the semantics of a program, such as changing a 00 to a $1$ or turning $x>y$ into $x<y$. The hypothesis is that a good test suite should be able to catch these artificial bugs, and there is a substantial evidence that the ability to catch both artificial and real-world bugs is strongly correlated (Just et al., 2014). In this paper, we make program edits that would not affect test cases. [^1] Reflection in TypeScript and Python allows a program to detect any source code change. But, most test suites do not use these features. Instead, we make syntactic edits that lead to type mispredictions for a given CodeLLM.
3 Methodology
3.1 Constructing Steering Datasets
Model choice
We build steering datasets for 1B and 7B parameter StarCoderBase models. These models are trained to fill-in-the-middle (§ 2), which is necessary for the type prediction task.
Source datasets
A distinguishing feature of our work is that we construct steering pairs from natural data (source code from GitHub) instead of using templates to construct simple examples. This provides more evidence that activation steering is robust. For Python, we use ManyTypes4Py (Mir et al., 2021), a dataset for neural type prediction. It features code from 5,382 Python projects that use Python type annotations and successfully type-check. For TypeScript, we filter The Stack (Kocetkov et al., 2023) to find 1.1M TypeScript files that type-check, of which we sample up to 2,000 for steering.
def __init__(self, x, y):
self.x = x
self.y = y
def delta_x(p: Point , x: float):
p.x = p.x + x
(a) The original program.
def __init__(self, x, y):
self.x = x
self.y = y
def delta_x(p: Type0, x: float):
p.x = p.x + x
(b) Type renaming.
def __init__(self, x, y):
self.x = x
self.y = y
def delta_x(p , x: float):
p.x = p.x + x
(d) Type annotation removal.
def __init__(self, x, y):
self.x = x
self.y = y
def delta_x(p: Point, tmp: float):
p.x = p.x + tmp
(c) Variable renaming.
Figure 2: Examples of three semantics-preserving edits. The type prediction site is float. We carefully ensure that each edit is internally consistent. E.g., in (2(c)), when we rename the binding x to tmp, we rename references to the binding. To construct a steering pair, we repeatedly apply edits until the CodeLLM mispredicts.
class KafkaAvroBackend(RepositoryBackend):
def __init__(
self, config__tmp0: dict, producer=AvroProducer, loader=AvroMessageLoader,
value_serializer: Callable = to_message_from_dto,
get_producer_config: Callable = get_producer_config,
get_loader_config: Callable = get_loader_config
) -> None:
producer_config = get_producer_config(config__tmp0)
loader_config = get_loader_config(config__tmp0)
Figure 3: A fragment of a Python steering pair. The original code is 70 lines of text. The dict is the expected prediction. But, renaming config to __tmp0 makes the model mispredict Repository, which is a hallucinated type.
Semantics-preserving code edits
For a given model $M$, our steering dataset is a set of triples $(x^{+},x^{-},t)$ where $t$ is a type and $x^{+},x^{-}$ are prompts that represent fill-in-the-middle type prediction tasks with $t$ as the expected type for both prompts. However, the maximum likelihood generation for $x^{+}$ is $M(x^{+})=t$ and for $x^{-}$ it is $M(x^{-})=k$, where $k\neq t$.
[…]
To construct the negative prompt, we incrementally apply semantics-preserving code edits to the positive prompt until $M$ mispredicts $t$. We implement the following edits. 1) Rename variable:We select a variable binding (from a function definition) and rename it to an arbitrary name that does not conflict with other variables in the program. We also rename all bound occurrences of the same variable so that the program’s semantics remain unchanged. 2) Remove type annotation:We select a type annotation (excluding the target $t$) and delete it. In a gradually typed language, removing or relaxing an annotation does not alter program semantics. 3) Rename user-defined type:We select an arbitrary type definition (e.g., a class name or a type alias) and rename it to an arbitrary name that does not conflict with other names in the program. 4) Rename builtin type:We also support renaming builtin types by introducing type aliases. Figure 2 illustrates several separate edits to a program.
By incrementally applying edits, the model eventually mispredicts $t$ (or else we fail to find an $x^{-}$ and thus discard $x^{+}$). Figure 3 illustrates a real example from our dataset where StarCoderBase-1B mispredicts after an edit. Note that a single edit often alters the positive prompt $x^{+}$ at several points. Moreover, when we apply several edits, the two prompts $x^{+},x^{-}$ may become significantly different from each other. We find that the mean edit-distance for our Python and TypeScript steering pairs is $74$ and $327$ respectively. This is in contrast to steering pairs constructed from sentence templates, that only differ by a small, fixed number of words at a far smaller number of places.
We hypothesize that mispredictions happen because the model fails to identify the inference task as a type prediction task. This likely occurs because our edits shift the prompt outside the distribution of a model’s training data. We thus call our edits distractors, since they distract from the correct task. We leverage activation steering to make the model more robust against such distractors and other semantically irrelevant features of code. For example, distractors often remove important in-context clues like function names which models rely on for prediction. Relying on textual meaning rather than program structure is a limitation of CodeLLMs that we improve with activation steering.
Class balance
The distribution of type annotations in natural data is heavily skewed: builtin types occur far more often than the long tail of user-defined types. To prevent bias towards majority-class types in our steering vectors, we balance every dataset $\mathcal{D}$ to ensure that no target type $t$ occurs more than $25$ times. In our evaluation, we also measure the effectiveness of steering by type (§ 4.2).
Python TypeScript
Edit Type Steering Held-out Steering Held-out
Rename variables $1,924$ $773$ $1,400$ $539$
Rename types $1,413$ $484$ $1,094$ $457$
Remove type annotations $1,952$ $892$ $698$ $297$
Rename variables & types $2,000$ $1,051$ $2,000$ $811$
Rename variables & remove type annotations $1,842$ $1,773$ $1,668$ $599$
Rename types & remove type annotations $2,000$ $781$ $1,619$ $445$
All edits $2,000$ $1,572$ $2,000$ $762$
Table 1: Overview of StarCoderBase-1B dataset sizes and applied code edits. Note that steering subsets are used for constructing steering vectors, while held out subsets are used to evaluate steering performance on out of distribution data. Dataset sizes vary because we only apply edits to programs where the model correctly predicts types. We then filter to remove overlap in source programs between held-out and steering sets. This accounts for the variation in sizes.
Ablation splits and evaluation sets
We construct several datasets $\mathcal{D}$, where each dataset applies a subset (or all) of the program edits described above to programs in a language. We further split each dataset into a steering subset and a held-out evaluation subset. We ensure that the two subsets have prompts based on different source programs, to test how steering generalizes to unseen programs. Table 1 summarizes the dataset sizes for StarCoderBase-1B. We report the sizes for StarCoderBase-7B in appendix A.
3.2 Constructing Steering Vectors
Given a dataset of steering pairs $(x_{i}^{+},x_{i}^{-},t)\in\mathcal{D}$ and model $M$, we apply a forward pass to every $M(x_{i}^{+}),M(x_{i}^{-})$ to collect model activations on the last token. Specifically, we extract activations from the residual stream of the model, which is the output of a transformer layer $\ell_{i}$ that in turn becomes the input to subsequent layers $\ell_{i+1...n}$. We write $A_{\ell}(x)$ to denote this activation tensor at some layer $\ell$ for prompt $x$. Thus we compute steering vectors $\mathbf{t}_{\ell}$—one for each layer—as the mean difference between positive and negative activations at that layer:
$\mathbf{t}_{\ell}=\frac{1}{|\mathcal{D}|}\sum_{(x_{i}^{+},x_{i}^{-},t)\in\mathcal{D}}\left(A_{\ell}(x_{i}^{+})-A_{\ell}(x_{i}^{-})\right)$ (1)
The intuition behind equation 1 is that the distance between positive and negative pairs in activation space encodes the transformation for steering towards the correct type.
[…]
4 Evaluation
4.1 Layer Ablations
Figure 4: Steering accuracy for StarCoderBase-1B on the held-out set of the TypeScript Datasets. Steering vectors are patched onto different subsets of adjacent layers.
We evaluate the accuracy of steering on each edit dataset (table 1). To find the optimal layer for applying steering vectors, we conduct an ablation on TypeScript data with StarCoderBase-1B. We consider patching at single layers as well as intervals of three and five adjacent layers. Figure 4 shows that for code edits where steering is most effective (e.g., renaming types), patching at single versus multiple layers makes little difference. However, for the less effective edits (e.g., removing type annotations), patching multiple layers significantly increases accuracy. We hypothesize that this performance gap exists because task vector refinement occurs over multiple layers, thus needs to be steered over multiple layers. This follows from previous ideas that transformers build predictions incrementally throughout layers (Elhage et al., 2021; Geva et al., 2022). Following this finding, we chose to patch layers 10–14 in StarCoderBase-1B and 19–23 in StarCoderBase-7B.
Python TypeScript
Edit Type Steering Random Steering Random
Rename variables 0.24 0.17 0.29 0.26
Rename types 0.90 0.10 0.75 0.11
Remove type annotations 0.48 0.20 0.41 0.31
Rename types & variables 0.56 0.13 0.49 0.09
Rename types & remove annotations 0.60 0.11 0.63 0.11
Rename variables & remove annotations 0.39 0.18 0.30 0.22
All edits 0.36 0.16 0.51 0.11
Python TypeScript
Edit Type Steering Random Steering Random
Rename variables 0.20 0.11 0.26 0.09
Rename types 0.84 0.07 0.69 0.07
Remove type annotations 0.49 0.11 0.39 0.18
Rename types & variables 0.56 0.10 0.36 0.05
Rename types & remove annotations 0.59 0.09 0.67 0.05
Rename variables & remove annotations 0.41 0.10 0.27 0.08
All edits 0.50 0.10 0.40 0.04
Table 2: Steering StarCoderBase-1B and 7B on several datasets of semantics-preserving code edits. For each dataset, we report the accuracy on a held out set of negative prompts, i.e., the model mispredicts all of these types without steering and steering corrects a significant number of mispredictions. The Random column reports accuracy after steering with a randomly initialized steering vector.
4.2 Evaluating Different Edits and Effectiveness By Type
Figure 5: Mean accuracy of steering Python by expected type label frequency. Error bars indicate the interquartile range. Note that we constrain the maximum size of any class to 25.
Having picked the layers to steer, table 2 summarizes the results of steering on each of our models. For StarCoderBase-1B, steering is most effective at fixing distractors from renaming types, achieving an accuracy of 90% in Python and 75% in TypeScript. We report accuracy over the held-out evaluation sets for each edit type. However, we also evaluated steering accuracy on the steering set and found no significant difference in performance between steering splits or held-out splits (appendix B). This is a first indication that our steering vectors generalize well.
To ensure that steering performance is not disproportionately affected by accuracy on a single type, we conduct an analysis of the results of steering per type label. In figure 5, we report steering accuracy on Python factored by target type frequency. Appendix C has a similar result for TypeScript. We find that the performance of steering vectors is well-distributed across types, and not skewed towards the majority class of labels.
4.3 Backup Circuits
A threat to validity for any intervention that involves activation patching is that the patch may not be directly improving performance, but just triggering a backup circuit (McGrath et al., 2023). For example, it could be that patching is just introducing noise into the embedding space that causes other mechanisms to trigger, thus producing the expected outcome. This effect complicates the interpretability of patches as well as steering vectors.
As part of our evaluation, we steer the model with a random steering vector (the Random columns in table 2). We find that steering in this way has non-zero accuracy, though it is generally significantly lower than our computed steering vectors. We posit that the non-zero accuracy of the random vector occurs due to backup circuits. However, the more significant gains from our computed steering vectors indicate that our steering method performs true transformations towards the correct type.
4.4 Steering Versus Finetuning
Fine-tuning is the traditional approach to improving model performance on a downstream task, so we now compare steering to fine-tuning. We fine-tune StarCoderBase-1B on the Python Renamed Types dataset. We fine-tune for five epochs using the AdamW optimizer and a learning rate of $3\times 10^{-5}$ with weight decay of $0.1$. At the end of each epoch, we evaluate on the held-out set.
In figure 6(a), we see that fine-tuning is approximately as effective as steering on the type prediction task. However, figure 6(b) shows that this comes with a caveat: fine-tuning on type-prediction significantly degrades the model’s code completion ability, as measured with HumanEval Chen et al. (2021). In contrast, since the steering vectors for type prediction are a small, localized patch, one can easily toggle them on or off based on the task at hand. On the other hand, fine-tuning for type prediction produces a new specialized model. We have shown that this fine-tuned model is weak at code completion, but in general, other abilities may have been impacted too.
(a) Type prediction accuracy.
(b) HumanEval pass@1.
Figure 6: We fine-tune StarCoderBase-1B on the Python Renamed Types dataset. We find that within 1 epoch, the fine-tuned model achieves the same accuracy on the held-out set as the steered model (figure 6(a)). However, this fine-tuning makes the model much worse at code completion (figure 6(b)).
4.5 Language Transfer in Steering Vectors
Edit Type $\mathrm{TS}\rightarrow\mathrm{Py}$ $\mathrm{Py}\rightarrow\mathrm{Py}$ $\mathrm{Py}\rightarrow\mathrm{TS}$ $\mathrm{TS}\rightarrow\mathrm{TS}$
Rename variables 0.24 0.24 0.20 0.29
Rename types 0.85 0.90 0.77 0.75
Remove type annotations 0.38 0.48 0.36 0.41
Rename types & variables 0.38 0.56 0.55 0.49
Rename types & remove annotations 0.60 0.60 0.64 0.63
Rename variables & remove annotations 0.26 0.39 0.36 0.30
All edits 0.27 0.36 0.34 0.51
Table 3: Accuracy of steering StarCoderBase-1B. A column labelled $A\rightarrow B$ indicates that the steering vector is computed from language $A$ but evaluated on language $B$.
Syntactically, Python and TypeScript look very different. However, semantically the two languages have a lot in common Politz et al. (2013); Bierman et al. (2014). In particular, both of them use gradual typing disciplines with many shared semantic features. So, we wondered, could CodeLLMs be learning a representation of types that is shared between these two languages?
To test this hypothesis, we test if steering vectors built on TypeScript data can improve the accuracy of Python type prediction, and vice versa. We conduct this experiment with each of our datasets: we steer StarCoderBase-1B using vectors from language $A$ but evaluate on the corresponding held-out test set from language $B$. Table 3 shows that this is nearly as effective as steering type prediction on the same language.
From this observation, we hypothesize that within the model there exists a representation of program types that is shared across TypeScript and Python. This is in contrast to the scenario where distinct model components are specialized to each programming language. Thus, steering vectors are effective across languages because they intervene on the shared representation between them. Moreover, these results suggest that steering vectors generalize by targeting a high-level representation of types.
5 Conclusion
We investigate activation steering for type prediction by making CodeLLMs more robust against semantically-irrelevant aspects of code. We find that by constructing steering pairs using semantics-preserving code edits we can construct highly effective steering vectors. Our experiments show that steering vectors generalize outside the programs used for steering and outperform naive baselines. We find that our steering vectors are transferable across languages. This suggests the existence of a shared representation of a type within the CodeLLM.
Activation steering can be a powerful technique for improving model performance on tasks where fine-tuning is not advantageous. As CodeLLMs are trained on more programming languages and downstream tasks, activation steering may be used as a lightweight alternative to multiple fine-tuned experts. Rather than fine-tuning several models on different tasks, the same base model could be specialized through steering vectors which are added and removed as needed. This could be especially convenient for resource-constrained applications. Our type prediction vectors, for example, could make a convenient lightweight expert for a coding assistant, useful for applications like type migration.
In future work, we aim to study the underlying causal mechanism in CodeLLMs responsible for type prediction. We further wish to explore how type-prediction steering vectors may generalize to open generation problems like code completion.
6 Ethical Considerations
The purpose of this work is to investigate methods for making CodeLLMs more robust and aligned to user intent. It is our view that interpreting CodeLLMs is necessary for understanding whether models approach programming in a principled way. As models become more integrated into developers’ workflows, model errors could compromise the security of entire systems. For this reason, we focus on making models more robust against distractors.
The datasets used in our investigation are sourced from publicly available code. Our TypeScript dataset is derived from a subset of The Stack v1.2, which contains permissively licensed data with personal identifying information (PII) filtered. The ManyTypes4Py dataset is funded by the European Commission, which follows data privacy laws under the EU General Data Protection Regulation (GDPR). We utilize the StarCoder family of models for their open and publicly available weights and training data.
7 Reproducibility
We commit to making all of our code and datasets public for the purpose of reproducibility and furthering research in the mechanistic interpretability of CodeLLMs.
Acknowledgments
We thank Ming-Ho Yee for providing the TypeScript dataset that we use in this work. We thank Northeastern Research Computing for providing the computing resources used for this work. This work is partially supported by the U.S. National Science Foundation (CCF-2052696).
Appendix A StarCoderBase-7B Steering Dataset Sizes
Table 4 reports the sizes of our steering and evaluation datasets for StarCoderBase-7B.
Python TypeScript
Edit Type Steering Held-out Steering Held-out
Rename variables $1,094$ 465 944 385
Rename types 796 343 773 332
Remove type annotations $1,528$ 706 875 344
Rename variables & types $1,622$ 759 $1,423$ 577
Rename variables & remove type annotations $2,000$ $1,246$ $1,270$ 478
Rename types & remove type annotations $2,000$ 977 893 383
All edits $2,000$ $1,044$ $1,663$ 630
Table 4: Overview of StarCoderBase-7B dataset sizes and applied code edits.
Appendix B Steering Accuracy on Construction Set
In § 4.2 we evaluated steering accuracy on held-out evaluation data. In this section, we evaluate accuracy on programs used to construct the steering vectors themselves. Our results in table 5 show that there is no significant difference in accuracy between held-out and steering sets.
Python TypeScript
Edit Type Steering Held-out Steering Held-out
Rename variables 0.23 0.24 0.29 0.29
Rename types 0.89 0.90 0.71 0.75
Remove type annotations 0.44 0.48 0.41 0.41
Rename types & variables 0.56 0.56 0.47 0.49
Rename types & remove annotations 0.60 0.60 0.61 0.63
Rename variables & remove annotations 0.39 0.39 0.33 0.30
All edits 0.42 0.36 0.49 0.51
Python TypeScript
Edit Type Steering Held-out Steering Held-out
Rename variables 0.19 0.20 0.29 0.26
Rename types 0.82 0.84 0.62 0.69
Remove type annotations 0.46 0.49 0.40 0.39
Rename types & variables 0.50 0.56 0.36 0.36
Rename types & remove annotations 0.52 0.59 0.63 0.67
Rename variables & remove annotations 0.33 0.41 0.29 0.27
All edits 0.46 0.50 0.40 0.40
Table 5: Steering StarCoderBase-1B and 7B on several datasets of semantics-preserving code edits. For each dataset, we report the accuracy on a held out set of negative prompts. Under the Steering column, we report the accuracy of steering vectors on the set of negative prompts used to construct the vectors.
Appendix C Accuracy by Type
We investigate whether steering vectors are biased towards particular types. Specifically, we look at whether steering vectors have higher accuracy on type labels that appear more frequently in the steering set. Typically, these types are builtin types such as string or str which are most commonly represented in type annotations. We look at the All Edits subsets of Python and Typescript in figure 5 and figure 7, respectively. For each type label frequency ranging from 1 to 25, we plot the mean accuracy of a steering vector on the held-out set. Results show that steering vectors are not biased towards more frequent type labels.
Figure 7: Accuracy by type for the TypeScript All Edits dataset. Error bars represent the interquartile range.

### v2

@@
-Activation Steering for Robust Type Prediction in CodeLLMs
+Understanding How CodeLLMs (Mis)Predict Types with Activation Steering
 Abstract
-Contemporary LLMs pretrained on code are capable of succeeding at a wide variety of programming tasks.
-However, their performance is very sensitive to syntactic features, such as the names of variables and types, the structure of code, and presence of type hints.
-We contribute an inference-time technique to make CodeLLMs more robust to syntactic distractors that are semantically irrelevant.
-Our methodology relies on activation steering, which involves editing internal model activations to steer the model towards the correct prediction.
-We contribute a novel way to construct steering vectors by taking inspiration from mutation testing, which constructs minimal semantics-breaking code edits.
-In contrast, we construct steering vectors from semantics-preserving code edits.
-We apply our approach to the task of type prediction for the gradually typed languages Python and TypeScript.
-This approach corrects up to 90% of type mispredictions.
-Finally, we show that steering vectors calculated from Python activations reliably correct type mispredictions in TypeScript, and vice versa.
-This result suggests that LLMs may be learning to transfer knowledge of types across programming languages.
+CodeLLMs are transforming software development as we know it.
+This is especially true for tasks where rule-based approaches fall short, like type prediction.
+The type prediction task consists in adding a new type annotation to a partially typed program, such that the resulting program is closer to being fully typed.
+The intractability of rule-based approaches and high cost of manual annotation make CodeLLMs an attractive solution to the problem.
+However, CodeLLMs are still far from being deployed on the large-scale due to doubts surrounding their reliability.
+To shed some light on how CodeLLMs approach type prediction, we investigate what happens when a model mispredicts a type.
+We show that by applying semantics-preserving edits to code, CodeLLMs are eventually misled into mispredicting type annotations.
+However, by leveraging activation steering we are able to “steer” the model back to the correct prediction, making models more robust against semantically irrelevant prompt features.
+We show that steering achieves comparable performance to fine-tuning directly on the type prediction task.
+Furthermore, we find that steering vectors computed from Python code are effective at correcting TypeScript mispredictions, and vice versa.
+To our knowledge, this is the first evidence of its kind to suggest that CodeLLMs learn task representations that transfer across languages.
+Understanding How CodeLLMs (Mis)Predict Types with Activation Steering
+Francesca Lucchetti & Arjun Guha Khoury College of Computer Sciences Northeastern University Boston, MA 02115 {lucchetti.f,a.guha}@northeastern.edu
 1 Introduction
-Large Language Models trained on code (CodeLLMs) are one of the most successful applications of LLMs to date.
-They are the foundation for both developer tools (e.g., Weiss & Yahav (2013); git (2021)) and for reasoning agents (e.g., Zelikman et al.
+Large language models trained on code, or CodeLLMs, are one of the most successful applications of LLMs to date.
+They are the foundation of both developer tools (e.g., Weiss and Yahav (2013); GitHub (2021)) and reasoning agents (e.g., Zelikman et al.
 (2024); Yang et al.
@@
 (2024); Li et al.
-(2022)).
-However, despite their utility, CodeLLMs can still be unreliable.
-Recent work shows that even the most capable models are not robust to small semantics-preserving changes such as refactorings and variable renamings (Hooda et al., 2024; Tambon et al., 2024).
-Thus CodeLLMs often cannot be deployed without careful supervision, such as a human-in-the-loop or robust code sandboxing.
-In this paper, we present an approach to make CodeLLMs more robust using activation steering (Li et al., 2024; Rimsky et al., 2023; Turner et al., 2023; Subramani et al., 2022).
-Activation steering is an inference-time model editing technique that modifies the intermediate computations (activations) of a model using steering vectors to steer model behavior toward desired outcomes.
-We present a novel technique for constructing steering vectors for the code domain by taking inspiration from mutation testing (DeMillo et al., 1978).
-Mutation testing automatically constructs minimal semantics-breaking code edits.
-In contrast, we construct steering vectors from minimal semantics-preserving code edits that lead to CodeLLM mispredictions.
-The nature of code allows us to construct these code edits in a sound and scaleable way.
+(2022)), capable of performing a variety of programming tasks across several programming languages.
+This paper focuses on type prediction for gradually typed programming languages, specifically Python and TypeScript.
+Given a partially typed program $p$ written in a gradually typed language, the type prediction task is defined as follows (Migeed and Palsberg, 2020):
 […]
-[…]
-Finally, we are surprised to find that type prediction steering transfers across programming languages.
-In particular, the steering vectors that we construct from TypeScript prompts are very effective at steering Python type predictions, and vice versa.
-We conduct a careful analysis and conclude that is it likely that multi-lingual CodeLLMs are learning a representation of types that is shared across the programming languages we study (§ 4.5).
-2 Background and Related Work
+A gradually typed programming language allows programs to freely mix typed and untyped code, giving programmers more flexibility than traditional static typing affords (Siek and Taha, 2006; Tobin-Hochstadt and Felleisen, 2006).
+However, this flexibility comes with a downside: at runtime, untyped code still needs to type-correct for the program to run correctly.
+With omitted or weak type annotations, type errors may not be caught until program execution.
+Automating type prediction is important for both scientific and practical reasons.
+From a practical perspective, well-known companies like Airbnb, Dropbox, Slack, Netflix, and many others have each taken several years to manually add type annotations to their multi-million line gradually typed codebases (Rudenko, 2020; Lehtosalo, 2019; Felix Rieseberg, 2017; Luke Autry, ; Sumana Mohan et al., 2022; Abacus, 2019; Mihai Parparita, 2020; Jake Zimmerman, 2022).
+Automated gradual type prediction would save significant effort and resources.
+From a scientific perspective, type prediction is intractable, and rule-based approaches have not been scaled up to contemporary languages (§ 2).
+Because of this, prior work has explored leveraging CodeLLMs for type prediction.
+However, CodeLLM performance on the task is far from being able to automate large-scale type annotation efforts.
+In particular, we show that CodeLLM type predictions are sensitive to small semantics-preserving changes in prompts like renaming variables, which limits their practical utility.
+The sensitivity of CodeLLMs towards trivial prompt changes begs the question of whether CodeLLMs approach type prediction in the “correct” way–that is, by inferring the type of a variable based on the operations performed over it.
+Alternatively, CodeLLMs could be relying on text features to make spurious predictions–for example, if a variable name is n, then its type is integer.
+We term these features “distractors”.
+As CodeLLM assistants are increasingly integrated into our workflows, understanding whether their predictions are principled or spurious is fundamental for security and large-scale adoption of these systems.
+To answer the question of how CodeLLMs approach type prediction, we design the following experimental setting.
+Given a type prediction prompt that the CodeLLM predicts correctly, we apply edits until the model output fails type checking.
+We construct our edits such that they do not matter for type prediction.
+To do so, we take inspiration from mutation testing (DeMillo et al., 1978).
+We present a variation of mutation testing that constructs minimal, semantics-preserving edits that trigger CodeLLM mispredictions.
+The nature of code allows us to construct these edits in a sound and scaleable way (§ 3.1).
+Since the model could originally predict the prompt, but fails with our trivial edits, this could mean that CodeLLM type prediction is spurious and based on distractors.
+However, our findings suggest a different story: that mispredictions happen because the model fails to identify the inference task as a type prediction task.
+This likely occurs because our edits shift the prompt outside the distribution of a model’s training data.
+We show that by editing intermediate computations (activations) of the model, we can recover performance on type prediction despite the presence of distractors.
+We employ activation steering (Li et al., 2024; Rimsky et al., 2023; Turner et al., 2023; Subramani et al., 2022) to steer model activations toward the target task of type prediction.
+Activation steering is an inference-time model editing technique that modifies the activations of a model using steering vectors to steer model behavior toward desired outcomes.
+Finally, we show that steering transfers across programming languages: the steering vectors we construct from TypeScript code are very effective at steering Python type predictions, and vice versa.
+This suggests that there exists a robust, generalizable mechanism for type prediction in CodeLLMs which, when activated, can correctly predict types despite the presence of distractors.
+From this, we conclude that CodeLLMs are learning task-solving mechanisms that transfer across languages (§ 4.4).
+To our knowledge, this is the first empirical evidence of its kind that identifies the occurrence of language transfer within LLMs  [^1] Transfer knowledge has previously been observed in models that were trained/fine-tuned on a single language but perform well on cross-lingual tasks.
+However, training data contamination challenges the validity of this result..
 return s[::-1] == s
@@
 Figure 1: An example of a code-infilling prompt using FIM special tokens. The correct prediction expected after the FIM middle token is str.
+2 Background and Related Work
+Classical type prediction and type inference
+Type prediction is also known as type migration, and is distinct from type inference as found in languages such as OCaml and Haskell.
+In those languages, every variable is typed, even if the types are implicit (Harper and Mitchell, 1993).
+In contrast, gradual type prediction can change program semantics (Phipps-Costin et al., 2021).
+There is work on gradual type prediction that uses human-written constraints and constraint solvers (Phipps-Costin et al., 2021; Rastogi et al., 2012; Siek and Vachharajani, 2008; Campora et al., 2018; Henglein and Rehof, 1995; Cartwright and Fagan, 1991).
+But, these papers present algorithms for variations of the lambda calculus or simple functional languages such as Scheme, and have not been scaled to more complex, modern programming languages.
 CodeLLMs and fill-in-the-middle
@@
 The aforementioned models are also trained to fill-in-the-middle (FIM) (Bavarian et al., 2022; Fried et al., 2023).
-FIM training does the following (1) It splits $\approx 50\%$ of training items into three chunks—prefix, middle, and suffix—of random lengths; (2) It adds special tokens to the start of each chunk to demarcate chunk boundaries; and (3) It reorders the middle chunk to appear last.
+FIM training does the following (1) It splits $\approx 50\%$ of training items into three chunks—prefix, middle, and suffix—of random lengths; (2) It adds special tokens to the start of each chunk; and (3) It reorders the middle chunk to appear last.
 The language modelling training objective remains unchanged.
@@
 Neural type prediction
-In this paper we study activation steering for type prediction, formulated as a fill-in-the-middle task.
-Prior work uses this formulation to evaluate base models (Yee & Guha, 2023; Fried et al., 2023), but does not study models’ internal mechanisms or try to improve task performance in a targeted way.
-There is prior work that trains smaller, specialized models for type prediction (Hellendoorn et al., 2018; Jesse et al., 2022; 2021; Pandi et al., 2021; Wei et al., 2020), but contemporary CodeLLMs outperform these specialized models (Yee & Guha, 2023; Fried et al., 2023).
-Classical type prediction and type inference
-Type prediction is also known as type migration, and is distinct from type inference as found in languages such as OCaml and Haskell.
-In those languages, every variable is typed, even if the types are implicit (Harper & Mitchell, 1993).
-In contrast, gradual type prediction can change program semantics (Phipps-Costin et al., 2021).
-There is work on gradual type prediction that uses human-written constraints and constraint solvers (Phipps-Costin et al., 2021; Rastogi et al., 2012; Siek & Vachharajani, 2008; Campora et al., 2018; Henglein & Rehof, 1995; Cartwright & Fagan, 1991).
-But, these papers present algorithms for variations of the lambda calculus or simple functional languages such as Scheme, and have not been scaled to more complex, modern programming languages.
-Mechanistic interpretability
-Interpretability seeks to understand the causal mechanisms behind model inference.
-In particular, mechanistic interpretability studies how model behaviors arise from specific components within the transformer architecture.
-One of the driving motivations of the field is that by understanding how a model works, we can understand how to align models to desired behavior.
-Prior work in mechanistic interpretability includes localizing and editing factual associations within the transformer (Meng et al., 2022), as well as probing hidden representations for knowledge of high-level concepts (Li et al., 2024; Dong et al., 2023).
-These works are forms of implicit evaluation of model ability, as opposed to explicit evaluation using benchmarks (Dong et al., 2023).
-Mechanistic interpretability has contributed a wealth of methods for implicit evaluation, for example activation patching (Vig et al., 2020; Variengien & Winsor, 2023), which applies patches to model activations in order to produce a change in behavior.
-Research leveraging activation patching has suggested the existence of task vectors (Hendel et al., 2023; Ilharco et al., 2022).
-Task vectors are representations within the transformer which encode a high-level notion of the task described by the prompt.
-The existence of task vectors is further supported by work in activation steering.
-Steering has been employed to diminish model deceitfulness and sycophancy (Rimsky et al., 2023; Li et al., 2024), which suggests that steering works by modifying the activations corresponding to task vectors.
-We thus draw on the idea of task vectors to perform our steering experiments.
-Importantly, steering relies on the linear representation hypothesis, which argues that concepts are represented as directions in the embedding space of a model (Park et al., 2023).
-We posit that successful steering performs transformations over model activations that shift the task representation towards a target direction.
-Despite much work on the mechanistic interpretability of LLMs, similar research on CodeLLMs has been limited.
-In this paper we focus on types, which are a fundamental feature of programming languages.
-One obstacle in this domain is that within a program, a type annotation is semantically constrained by both the tokens that precede it and those that succeed it.
+We follow prior work in formulating type prediction as a fill-in-the-middle task  (Yee and Guha, 2023; Fried et al., 2023).
+Within a program, a type annotation is semantically constrained by both the tokens that precede it and those that succeed it.
 For example, the type of a function argument is constrained by how the argument is subsequently used in the function body.
-To overcome this, we use CodeLLMs that are trained on FIM.
-This allows us to construct prompts for type prediction.
+FIM allows us to construct prompts for type prediction as next token prediction, moving both function body and signature to the prompt prefix.
+Prior work also includes finetuning smaller, specialized models for type prediction (Hellendoorn et al., 2018; Jesse et al., 2022, 2021; Pandi et al., 2021; Wei et al., 2020), but contemporary CodeLLMs outperform them.
+It is now well known that even the most capable CodeLLMs are sensitive to small variations in prompts (Hooda et al., 2024; Tambon et al., 2024).
+Prior work uses a black-box approach to study this phenomena.
+In contrast, we investigate type prediction in depth using a whitebox approach, to shed light on the causes of this sensitivity.
+We further investigate editing internal model components to improve models’ robustness against prompt variations.
 Mutation testing and program transformations
-We take inspiration from mutation testing (DeMillo et al., 1978), but our technique is different.
+To build our semantics-preserving edits, we take inspiration from mutation testing (DeMillo et al., 1978).
 The goal of mutation testing is to test a program’s test suite.
@@
 The hypothesis is that a good test suite should be able to catch these artificial bugs, and there is a substantial evidence that the ability to catch both artificial and real-world bugs is strongly correlated (Just et al., 2014).
-In this paper, we make program edits that would not affect test cases.
+Our technique differs from mutation testing in one key way: we make program edits that would not affect test cases.
 [^2] Reflection in TypeScript and Python allows a program to detect any source code change.
@@
 Instead, we make syntactic edits that lead to type mispredictions for a given CodeLLM.
+Mechanistic interpretability
+Mechanistic interpretability has contributed a wealth of methods for investigating models’ internal mechanisms, for example activation patching (Vig et al., 2020; Variengien and Winsor, 2023).
+Recently, research leveraging activation patching has suggested the existence of task vectors (Hendel et al., 2023; Ilharco et al., 2022)–latent representations within the transformer which encode a high-level notion of the task described in the prompt.
+The existence of task vectors is further supported by work in activation steering.
+Steering has been employed to control qualities like deceitfulness and sycophancy in model outputs (Rimsky et al., 2023; Li et al., 2024), which suggests that steering works by modifying the activations corresponding to task vectors.
+The connection between task vectors and steering is what motivates our investigation into steering for understanding type prediction.
+Despite much work on the mechanistic interpretability of LLMs, similar research on CodeLLMs has been limited, and research using real-world (e.g.
+GitHub) code is even more so.
+This is because MI techniques are hard to adapt across different prompt formats Hobbhahn et al.
+(2022).
+We are able to overcome this by using FIM tokens to standardize our patching experiments.
 3 Methodology
@@
 Model choice
-We build steering datasets for 1B and 7B parameter StarCoderBase models. These models are trained to fill-in-the-middle (§ 2), which is necessary for the type prediction task.
+We build steering datasets for 1B and 7B parameter StarCoderBase models Li et al. (2023). These models are both trained to fill-in-the-middle (§ 2).
 Source datasets
-A distinguishing feature of our work is that we construct steering pairs from natural data (source code from GitHub) instead of using templates to construct simple examples.
-This provides more evidence that activation steering is robust.
+A distinguishing feature of our work is that we construct steering pairs from natural data (source code from GitHub) instead of using prompt templates to construct simple examples.
+This is to ensure that activation steering is robust across varied examples.
 For Python, we use ManyTypes4Py (Mir et al., 2021), a dataset for neural type prediction.
@@
 The dict is the expected prediction.
-But, renaming config to __tmp0 makes the model mispredict Repository, which is a hallucinated type.
+But, renaming config to __tmp0 makes the model mispredict Repository, which is a hallucination.
 Semantics-preserving code edits
-For a given model $M$, our steering dataset is a set of triples $(x^{+},x^{-},t)$ where $t$ is a type and $x^{+},x^{-}$ are prompts that represent fill-in-the-middle type prediction tasks with $t$ as the expected type for both prompts.
-However, the maximum likelihood generation for $x^{+}$ is $M(x^{+})=t$ and for $x^{-}$ it is $M(x^{-})=k$, where $k\neq t$.
+For a given model $M$, our steering dataset is a set of triples $(x^{+},x^{-},t)$ where $x^{+},x^{-}$ are FIM type prediction prompts and $t$ is the expected type for both prompts.
+However, the maximum likelihood generation for $x^{+}$ is $M(x^{+})=t$ and for $x^{-}$ it is $M(x^{-})=t^{\prime}$, where $t^{\prime}\neq t$.
 […]
@@
 We implement the following edits.
-1) Rename variable:We select a variable binding (from a function definition) and rename it to an arbitrary name that does not conflict with other variables in the program.
+1) Rename variable:We select a variable binding (from a function definition) and rename it to an arbitrary name that does not conflict with other variables.
 We also rename all bound occurrences of the same variable so that the program’s semantics remain unchanged.
@@
 We find that the mean edit-distance for our Python and TypeScript steering pairs is $74$ and $327$ respectively.
-This is in contrast to steering pairs constructed from sentence templates, that only differ by a small, fixed number of words at a far smaller number of places.
-We hypothesize that mispredictions happen because the model fails to identify the inference task as a type prediction task.
-This likely occurs because our edits shift the prompt outside the distribution of a model’s training data.
-We thus call our edits distractors, since they distract from the correct task.
-We leverage activation steering to make the model more robust against such distractors and other semantically irrelevant features of code.
-For example, distractors often remove important in-context clues like function names which models rely on for prediction.
-Relying on textual meaning rather than program structure is a limitation of CodeLLMs that we improve with activation steering.
-Class balance
-The distribution of type annotations in natural data is heavily skewed: builtin types occur far more often than the long tail of user-defined types.
-To prevent bias towards majority-class types in our steering vectors, we balance every dataset $\mathcal{D}$ to ensure that no target type $t$ occurs more than $25$ times.
-In our evaluation, we also measure the effectiveness of steering by type (§ 4.2).
 Python TypeScript
@@
 Note that steering subsets are used for constructing steering vectors, while held out subsets are used to evaluate steering performance on out of distribution data.
-Dataset sizes vary because we only apply edits to programs where the model correctly predicts types.
+Dataset sizes vary because we only apply edits to programs for which the model predicts correct types.
 We then filter to remove overlap in source programs between held-out and steering sets.
-This accounts for the variation in sizes.
+Class balance
+The distribution of type annotations in natural data is heavily skewed: builtin types occur far more often than the long tail of user-defined types.
+To prevent bias towards majority-class types in our steering vectors, we balance every dataset $\mathcal{D}$ to ensure that no target type $t$ occurs more than $25$ times.
+In our evaluation, we also measure the effectiveness of steering by type (§ 4.2).
 Ablation splits and evaluation sets
-We construct several datasets $\mathcal{D}$, where each dataset applies a subset (or all) of the program edits described above to programs in a language.
+We construct several datasets $\mathcal{D}$, where each dataset applies a subset (or all) of the aforementioned edits to programs in a language.
 We further split each dataset into a steering subset and a held-out evaluation subset.
-We ensure that the two subsets have prompts based on different source programs, to test how steering generalizes to unseen programs.
+We ensure that the two subsets have prompts based on different source programs.
+This tests how well steering generalizes to wholly unseen programs.
 Table 1 summarizes the dataset sizes for StarCoderBase-1B.
@@
 3.2 Constructing Steering Vectors
-Given a dataset of steering pairs $(x_{i}^{+},x_{i}^{-},t)\in\mathcal{D}$ and model $M$, we apply a forward pass to every $M(x_{i}^{+}),M(x_{i}^{-})$ to collect model activations on the last token.
+Given a dataset of steering pairs $(x_{i}^{+},x_{i}^{-},t)\in\mathcal{D}$ and model $M$, we apply forward passes $M(x_{i}^{+}),M(x_{i}^{-})$ and save model activations on the last token.
 Specifically, we extract activations from the residual stream of the model, which is the output of a transformer layer $\ell_{i}$ that in turn becomes the input to subsequent layers $\ell_{i+1...n}$.
@@
 $\mathbf{t}_{\ell}=\frac{1}{|\mathcal{D}|}\sum_{(x_{i}^{+},x_{i}^{-},t)\in\mathcal{D}}\left(A_{\ell}(x_{i}^{+})-A_{\ell}(x_{i}^{-})\right)$ (1)
-The intuition behind equation 1 is that the distance between positive and negative pairs in activation space encodes the transformation for steering towards the correct type.
+The intuition behind equation 1 is that the distance between positive and negative pairs in activation space encodes the transformation for steering towards the correct type prediction.
 […]
-4 Evaluation
-4.1 Layer Ablations
+In the next section, we perform steering at the level of both individual layers and sets of adjacent layers simultaneously.
+4 Results
+In this section, we evaluate model performance on type prediction after activation steering. Note that performance before steering is 0.
+4.1 At What Layers is Steering Most Effective?
 Figure 4: Steering accuracy for StarCoderBase-1B on the held-out set of the TypeScript Datasets. Steering vectors are patched onto different subsets of adjacent layers.
-We evaluate the accuracy of steering on each edit dataset (table 1).
-To find the optimal layer for applying steering vectors, we conduct an ablation on TypeScript data with StarCoderBase-1B.
-We consider patching at single layers as well as intervals of three and five adjacent layers.
-Figure 4 shows that for code edits where steering is most effective (e.g., renaming types), patching at single versus multiple layers makes little difference.
+To find the optimal layer for activation steering, we conduct a layer ablation using TypeScript data with StarCoderBase-1B (figure 4).
+We evaluate the accuracy of steering on each edit dataset (table 1), and we explore patching at single layers as well as intervals of three and five adjacent layers.
+There are two takeaways from this experiment.
+First, we find that steering middle and later layers is most effective.
+In the early layers of the model, the hidden state of the last token represents the last token itself.
+As computation moves through layers and layer outputs are incrementally added to the residual stream, this hidden state is refined into the next token prediction.
+This phase change in computation coincides with the model identifying the inference task, thus makes the target of our steering.
+Secondly, we show that steering at multiple layers improves performance.
+Specifically, for code edits where steering is most effective (e.g., renaming types), patching at single versus multiple layers makes little difference.
 However, for the less effective edits (e.g., removing type annotations), patching multiple layers significantly increases accuracy.
-We hypothesize that this performance gap exists because task vector refinement occurs over multiple layers, thus needs to be steered over multiple layers.
+We hypothesize that this performance gap exists because task vector refinement occurs over multiple layers.
+In these cases, activations need to be steered over multiple layers.
 This follows from previous ideas that transformers build predictions incrementally throughout layers (Elhage et al., 2021; Geva et al., 2022).
-Following this finding, we chose to patch layers 10–14 in StarCoderBase-1B and 19–23 in StarCoderBase-7B.
 Python TypeScript
@@
 All edits 0.50 0.10 0.40 0.04
-Table 2: Steering StarCoderBase-1B and 7B on several datasets of semantics-preserving code edits.
-For each dataset, we report the accuracy on a held out set of negative prompts, i.e., the model mispredicts all of these types without steering and steering corrects a significant number of mispredictions.
+Table 2: Steering StarCoderBase-1B and 7B on datasets of semantics-preserving edits.
+For each dataset, we report the accuracy on a held out set of negative prompts, i.e., the model mispredicts all of these types before steering.
 The Random column reports accuracy after steering with a randomly initialized steering vector.
-4.2 Evaluating Different Edits and Effectiveness By Type
-Figure 5: Mean accuracy of steering Python by expected type label frequency. Error bars indicate the interquartile range. Note that we constrain the maximum size of any class to 25.
-Having picked the layers to steer, table 2 summarizes the results of steering on each of our models.
-For StarCoderBase-1B, steering is most effective at fixing distractors from renaming types, achieving an accuracy of 90% in Python and 75% in TypeScript.
-We report accuracy over the held-out evaluation sets for each edit type.
-However, we also evaluated steering accuracy on the steering set and found no significant difference in performance between steering splits or held-out splits (appendix B).
+4.2 Does Steering Work Across Types and Different Kinds of Edits?
+We now consider the effectiveness of steering for different kinds of edits.
+Based on the results of the previous section, we do so with patches applied to layers 10–14 for StarCoderBase-1B and 19–23 for StarCoderBase-7B.
+Table 2 summarizes the results under the Steering columns.
+We report accuracy over held-out evaluation sets for each edit type.
+The effectiveness of steering can vary significantly.
+For StarCoderBase-1B, steering is most effective at fixing distractors from renaming types, achieving an accuracy of 90% for Python and 75% for TypeScript.
+(Note that without steering, the models mispredict on all problems in the dataset.) We also evaluate steering accuracy on the steering set and find no significant difference in performance between steering splits or held-out splits (appendix B).
 This is a first indication that our steering vectors generalize well.
@@
 We find that the performance of steering vectors is well-distributed across types, and not skewed towards the majority class of labels.
-4.3 Backup Circuits
+Figure 5: Mean accuracy of steering Python by expected type label frequency. Error bars indicate the interquartile range. Each class has maximum size of 25.
+4.3 How Does Steering Compare to Baselines?
+Backup Circuits
 A threat to validity for any intervention that involves activation patching is that the patch may not be directly improving performance, but just triggering a backup circuit (McGrath et al., 2023).
 For example, it could be that patching is just introducing noise into the embedding space that causes other mechanisms to trigger, thus producing the expected outcome.
-This effect complicates the interpretability of patches as well as steering vectors.
-As part of our evaluation, we steer the model with a random steering vector (the Random columns in table 2).
-We find that steering in this way has non-zero accuracy, though it is generally significantly lower than our computed steering vectors.
-We posit that the non-zero accuracy of the random vector occurs due to backup circuits.
-However, the more significant gains from our computed steering vectors indicate that our steering method performs true transformations towards the correct type.
-4.4 Steering Versus Finetuning
-Fine-tuning is the traditional approach to improving model performance on a downstream task, so we now compare steering to fine-tuning.
-We fine-tune StarCoderBase-1B on the Python Renamed Types dataset.
-We fine-tune for five epochs using the AdamW optimizer and a learning rate of $3\times 10^{-5}$ with weight decay of $0.1$.
-At the end of each epoch, we evaluate on the held-out set.
-In figure 6(a), we see that fine-tuning is approximately as effective as steering on the type prediction task.
-However, figure 6(b) shows that this comes with a caveat: fine-tuning on type-prediction significantly degrades the model’s code completion ability, as measured with HumanEval Chen et al.
-(2021).
-In contrast, since the steering vectors for type prediction are a small, localized patch, one can easily toggle them on or off based on the task at hand.
-On the other hand, fine-tuning for type prediction produces a new specialized model.
-We have shown that this fine-tuned model is weak at code completion, but in general, other abilities may have been impacted too.
-(a) Type prediction accuracy.
-(b) HumanEval pass@1.
-Figure 6: We fine-tune StarCoderBase-1B on the Python Renamed Types dataset.
-We find that within 1 epoch, the fine-tuned model achieves the same accuracy on the held-out set as the steered model (figure 6(a)).
-However, this fine-tuning makes the model much worse at code completion (figure 6(b)).
-4.5 Language Transfer in Steering Vectors
+This effect complicates the interpretability of steering vectors.
+For a baseline, we steer the model with a random steering vector (the Random columns in table 2).
+We find that steering this way has non-zero accuracy, but it is significantly lower than our computed steering vectors.
+Thus we conclude that the steering vectors we compute perform true, localized transformations towards the correct type prediction task in activation space.
 Edit Type $\mathrm{TS}\rightarrow\mathrm{Py}$ $\mathrm{Py}\rightarrow\mathrm{Py}$ $\mathrm{Py}\rightarrow\mathrm{TS}$ $\mathrm{TS}\rightarrow\mathrm{TS}$
@@
 Table 3: Accuracy of steering StarCoderBase-1B. A column labelled $A\rightarrow B$ indicates that the steering vector is computed from language $A$ but evaluated on language $B$.
-Syntactically, Python and TypeScript look very different.
-However, semantically the two languages have a lot in common Politz et al.
+Fine-tuning
+It is not yet clear if the accuracy of our steering is “good”.
+To answer this question, we compare steering to fine-tuning.
+Intuitively, fine-tuning a model on a narrow task gives us an upper bound on how good the model can be at that task.
+We fine-tune StarCoderBase-1B on the Python Renamed Types dataset.
+We fine-tune for five epochs using the AdamW optimizer and a learning rate of $3\times 10^{-5}$ with weight decay of $0.1$.
+At the end of each epoch, we evaluate on the held-out set.
+We see that fine-tuning is approximately as effective as steering (figure 6).
+This further indicates that steering targets and steers the part of the model responsible for the task.
+Figure 6: We fine-tune StarCoderBase-1B on the Python Renamed Types dataset. We find that within 1 epoch, the fine-tuned model achieves the same accuracy on the held-out set as the steered model.
+4.4 Do Models Learn Type Prediction Across Languages?
+Python and TypeScript are syntactically distinct, but their semantics have a lot in common Politz et al.
 (2013); Bierman et al.
 (2014).
-In particular, both of them use gradual typing disciplines with many shared semantic features.
-So, we wondered, could CodeLLMs be learning a representation of types that is shared between these two languages?
-To test this hypothesis, we test if steering vectors built on TypeScript data can improve the accuracy of Python type prediction, and vice versa.
+Both languages use gradual typing disciplines with many shared semantic features.
+So, could CodeLLMs be learning a mechanism for type prediction that is language-agnostic?
+To test this hypothesis, we evaluate if steering vectors built on TypeScript data can improve the accuracy of Python type prediction, and vice versa.
 We conduct this experiment with each of our datasets: we steer StarCoderBase-1B using vectors from language $A$ but evaluate on the corresponding held-out test set from language $B$.
 Table 3 shows that this is nearly as effective as steering type prediction on the same language.
-From this observation, we hypothesize that within the model there exists a representation of program types that is shared across TypeScript and Python.
-This is in contrast to the scenario where distinct model components are specialized to each programming language.
-Thus, steering vectors are effective across languages because they intervene on the shared representation between them.
-Moreover, these results suggest that steering vectors generalize by targeting a high-level representation of types.
+Collectively, results indicate that our steering vectors do not represent a specific type token like str or int.
+Instead, they encode a delta which steers model generations toward a mechanism for type prediction that 1) generalizes across different source codes; 2) is less sensitive to semantically irrelevant distractors; 3) generalizes across the languages we study.
+Given these observations, we conclude that there exists a non-spurious mechanism for type prediction in CodeLLMs which, when activated either natively or through activation steering, is more robust against distractors.
 5 Conclusion
-We investigate activation steering for type prediction by making CodeLLMs more robust against semantically-irrelevant aspects of code.
-We find that by constructing steering pairs using semantics-preserving code edits we can construct highly effective steering vectors.
-Our experiments show that steering vectors generalize outside the programs used for steering and outperform naive baselines.
-We find that our steering vectors are transferable across languages.
-This suggests the existence of a shared representation of a type within the CodeLLM.
-Activation steering can be a powerful technique for improving model performance on tasks where fine-tuning is not advantageous.
-As CodeLLMs are trained on more programming languages and downstream tasks, activation steering may be used as a lightweight alternative to multiple fine-tuned experts.
-Rather than fine-tuning several models on different tasks, the same base model could be specialized through steering vectors which are added and removed as needed.
-This could be especially convenient for resource-constrained applications.
-Our type prediction vectors, for example, could make a convenient lightweight expert for a coding assistant, useful for applications like type migration.
-In future work, we aim to study the underlying causal mechanism in CodeLLMs responsible for type prediction.
-We further wish to explore how type-prediction steering vectors may generalize to open generation problems like code completion.
-6 Ethical Considerations
-The purpose of this work is to investigate methods for making CodeLLMs more robust and aligned to user intent.
+We investigate how CodeLLMs perform gradual type prediction using the two most widely used gradually typed languages: TypeScript and Python.
+We leverage activation steering and FIM prompting as interpretability tools to investigate what happens when a model mispredicts.
+By using semantics-preserving code edits, we construct distractors that cause the model to mispredict.
+We then show that we can recover performance by constructing steering vectors to guide the model towards more robust type prediction.
+Our steering vectors are as effective as fine-tuning a model on type prediction.
+This indicates that our vectors effectively steer the part of the model responsible for type prediction.
+Furthermore, Python steering vectors are as effective at steering TypeScript, and vice versa.
+This provides new evidence that type prediction in CodeLLMs is not spurious, but a causal mechanism learned across languages that generalizes beyond syntax.
+6 Limitations
+Our findings shed light on how CodeLLMs perform the type prediction task for TypeScript and Python.
+Both these languages are well represented in CodeLLM training corpora.
+However, our findings may not generalize to low-resource gradually typed languages, e.g., Typed Racket Tobin-Hochstadt and Felleisen (2008) or Luau Lily Brown et al.
+(2023) since the performance of base models on these languages is very poor.
+We show that our approach works for two sizes of the StarCoderBase model family.
+We chose these models because we found them to be more effective at type prediction than other models trained with FIM Since StarCoder training data is fully open, we know the quantity and quality of StarCoder FIM training is high–a necessary requirement for fairly evaluating steering on FIM type prediction.
+7 Ethics Statement
+The purpose of this work is to understand whether CodeLLMs perform type prediction using reliable mechanisms.
 It is our view that interpreting CodeLLMs is necessary for understanding whether models approach programming in a principled way.
-As models become more integrated into developers’ workflows, model errors could compromise the security of entire systems.
+As CodeLLMs become more integrated into developers’ workflows, model errors could compromise the security of entire systems.
 For this reason, we focus on making models more robust against distractors.
-The datasets used in our investigation are sourced from publicly available code.
+The StarCoderBase models have an OpenRAIL license which allows the kind of investigation in this paper.
 Our TypeScript dataset is derived from a subset of The Stack v1.2, which contains permissively licensed data with personal identifying information (PII) filtered.
 The ManyTypes4Py dataset is funded by the European Commission, which follows data privacy laws under the EU General Data Protection Regulation (GDPR).
+These datasets are intended for CodeLLMs, which this paper investigates.
 We utilize the StarCoder family of models for their open and publicly available weights and training data.
-7 Reproducibility
-We commit to making all of our code and datasets public for the purpose of reproducibility and furthering research in the mechanistic interpretability of CodeLLMs.
+The size and efficient implementation of attention in these models allows us to use limited compute resources.
+We were able to run experiments on a single A100-80GB GPU.
+In total, we estimate around 48h of GPU time, including preliminary experiments and fine-tuning.
+Note that the computation of our final steering vectors is a small fraction of this time.
 Acknowledgments
-We thank Ming-Ho Yee for providing the TypeScript dataset that we use in this work.
+Portions of this work are implemented with NNSight and NDIF Fiotto-Kaufman et al.
+(2024).
+We thank Ming-Ho Yee for help with the TypeScript dataset that we use in this work Yee (2024).
 We thank Northeastern Research Computing for providing the computing resources used for this work.
-This work is partially supported by the U.S.
-National Science Foundation (CCF-2052696).
 Appendix A StarCoderBase-7B Steering Dataset Sizes